UT 20130914 AO#2 DX

Observer:JCGuerra(LBTO),SEsposito,LBuzoni,APuglisi,RBruglio (Arcetri Observatorio)
Telescope Operator:
SW Support:
AO Support:
Telescope Support:
Instrument Support: (LBTO)


  • Py-WFS working without any problem.
  • Installation of the mouse hole mask.
  • Acquisition of first intmat to see signal in the WFS and the pupils..
  • 4D for flattening the mirror.
  • Alignment of the WFS, rotation and translation of the subapertures vs ASM actuators.
  • Acquisition of new pupils
  • Acquisition of INTMAT to test with the mask.
  • Acquisition of IRTC psf.
  • A first interesting and good results with 400 modes.
  • Check of the mouse hole cover after the RR was removed
  • Cloudy and high humidity during the twilight
  • Hand-over 5pm
  • Arcetri report for detailed info


The mask is modified with three holes to screw the three teflon screws.

The mouse hole mask is installed in a configuration of ASM retracted to have access from the front. The mask is inserted easily without any tight resistance. The mask is tight up and the three screws are hand tide, the mask seems well fitted.

The acquisition of the intmat with the mask hole which block the first ring of actuators. The reconstructor built didn't work and we can see the inner ring is too high. Another problem has been identified at teh edge of the mirror where the WFS is not seeing. It is discussed to run the 4D to get a better flat of the edges.

A new flat is measured and saved, this new flat shows a better flat at the edges.

track number # modes WF rms
20130914_170324 660 70nm

Move M3 for the WFS focal station and a first look of the pupils show an uniform shape, this is the starting point for the alignment and acquisitions of pupils and Intmat.

It is decided to digital obscure a few subaperture in the centre as a first try, after looking the interferogram. The acquisition of the intmat is done with 400 modes applied and the test of the reconstructor at close loop is not good, the effect in the centre is too high and the close loop breaks.

After looking the data is decided to reduce the digital obscuration to get a configuration of 4x3 pixels for the central obscuration. This configuration match the on sky central obscuration.

With this new pupil shapes, the new acquisition and recorded of pupils is done and used for the intmat . We started to align the WFS in translation and rotation using the vertical and horizontal patterns. This requires many iteration of this process and at the end we are having an alignment inside of 0.1 pixels

The WFS is aligned with the new pupils and the acquisiton of new INTMAT started.

After some tries with different reconstructors built , we have reached to get one that it is working stable with 400 modes. This reconstructor didn't show any features in the inner rings and it responds very well, the loop was close for about 10 minutes and it didn't show any problem. The system is running at close loop with the local atmopshere, that means that we are not injecting any disturbance to the secondary. With this reconstructor a few track number were acquired with the IRTC PSF. The result is looking promising. Strehl of the PSF without disturbance about 85% in H band

Resume of the data collected:

TrackNo RefStar Mag El Wind DIMM/OL Rec bin #mod freq gain mod nph AntiDrift SR filter exp #frames disturb SN notes
20130914_234631 ??? 6.6 90 5 NaN -1.00 232352 1 400 990 0.6 0.2 0.4 3 1427 ON 84.5 H 20 500 OFF zero1600.fits  
20130914_234724 ??? 6.6 90 4 NaN -1.00 232352 1 400 990 0.6 0.2 0.4 3 1406 ON 85.8 H 20 500 OFF zero1600.fits  
20130914_234959 ??? 6.6 90 4 NaN -1.00 232352 1 400 990 0.6 0.2 0.4 3 1432 ON 84.7 H 20 500 OFF zero1600.fits  

-- JuanGuerra - 15 Sep 2013
  • psf_bin1_400modes:
Topic revision: r4 - 21 Sep 2013, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback